# Adding Multi-modal Controls to Whole-body Human Motion Generation

### v1

@@ Abstract
Whole-body multi-modal motion generation, controlled by text, speech, or music, has numerous applications including video generation and character animation. However, employing a unified model to accomplish various generation tasks with different condition modalities presents two main challenges: motion distribution drifts across different generation scenarios and the complex optimization of mixed conditions with varying granularity. Furthermore, inconsistent motion formats in existing datasets further hinder effective multi-modal motion generation. In this paper, we propose ControlMM, a unified framework to Control whole-body Multi-modal Motion generation in a plug-and-play manner. To effectively learn and transfer motion knowledge across different motion distributions, we propose ControlMM-Attn, for parallel modeling of static and dynamic human topology graphs. To handle conditions with varying granularity, ControlMM employs a coarse-to-fine training strategy, including stage-1 text-to-motion pre-training for semantic generation and stage-2 multi-modal control adaptation for conditions of varying low-level granularity. To address existing benchmarks’ varying motion format limitations, we introduce ControlMM-Bench, the first publicly available multi-modal whole-body human motion generation benchmark based on the unified whole-body SMPL-X format. Extensive experiments show that ControlMM achieves state-of-the-art performance across various standard motion generation tasks. Our website is at https://yxbian23.github.io/ControlMM. Corresponding authors.
Figure 1: We present ControlMM, a whole-body motion generation framework with plug-and-play multi-modal control, which can generalize robust motion generation abilities across Text-to-Motion, Speech-to-Gesture, and Music-to-Dance.
@@ Introduction
Whole-body human motion generation with multi-modal control  (Zhang et al. 2024b; Liu et al. 2024a; Li et al. 2023), which produces natural and logical human movements based on multi-modal conditions, has many downstream applications, including human video generation (Hu 2024) and character animation (Zhang et al. 2023a).
Recent advancements in single-conditioned human motion generation have enabled realistic human motion generation under control signals from varying modalities, including text descriptions (Guo et al. 2022; Zhang et al. 2023c), music clips (Siyao et al. 2022; Li et al. 2023), and speech segments (Liu et al. 2024a; Chen et al. 2024). However, achieving whole-body motion generation with multi-modal control in a unified model encounters several serious challenges: ❶ Motion distribution drift: The motion distribution in generated tasks corresponding to different control signals exhibits significant distribution drift. Under limited available motion data compared with video/text, current methods struggle to transfer robust motion topology knowledge across different multi-modal motion generation scenarios (Ling et al. 2023; Zhang et al. 2024b). ❷ Struggling optimization under mixed conditions: Current multi-modal motion generation methods mainly choose to compress control signals of varying granularity into a shared latent space to learn the multi-modal conditioned generation abilities (Zhou, Wan, and Wang 2023; Zhang et al. 2024b), leading to alignment issues across different modalities and optimization confusion when learning conditions at different granularity simultaneously. ❸ Non-uniform motion format and evaluation: There are no multi-modal whole-body human motion generation benchmarks with unified motion format and evaluation pipelines, impeding this field’s progress.
@@
  • Struggling optimization under mixed conditions: Current multi-modal motion generation work often compresses different control signals (semantic text guidance, low-level speeches on first-person perspective and the low-level music on third-person perspective) into a latent space for mixed modeling, including transformer token embedding (Zhou, Wan, and Wang 2023) and ImageBind’s (Girdhar et al. 2023) hidden space (Zhang et al. 2024b). However, due to the significant information gaps between different modalities, such compression and mixed learning leads to notable optimization burden (Team et al. 2023). Therefore, we choose to introduce different control signals through separate control branches and conduct a coarse-to-fine progressive training strategy to cultivate generation abilities under conditions of varying granularity. This also provides our model with plug-and-play, user-friendly, and efficient training properties.
  • Non-uniform motion format and evaluation: Finally, there are no high-quality multi-modal whole-body human motion generation benchmarks with unified motion representation and evaluation pipelines, which prompted us to build the first such benchmarks based on the SMPL-X (Pavlakos et al. 2019) whole-body format.
@@
To address existing benchmarks’ varying motion format limitations, we also built ControlMM-Bench, the first publicly available multi-modal human motion generation benchmark based on the unified whole-body representation SMPL-X, including unified data construction and evaluation pipelines. Extensive experiments show that ControlMM achieves competitive performance across various standard motion generation tasks, including text-to-motion, speech-to-gesture, and music-to-dance, compared to state-of-the-art baselines. We also provide extensive ablation studies offering valuable insights into model design, training strategies, and scaling effects for future multi-modal whole-body human motion generation foundation models.
@@
  • In this work, we decompose the motion distribution drift modeling in multi-modal generation into the static human skeleton structure capture and dynamic topology relationship measurement. For the first time, our ControlMM-Attn achieves parallel modeling of such human-centric spatial properties to boost whole-body motion generation against distribution drifts under varying conditions.
  • We introduce ControlMM, a two-stage coarse-to-fine multi-modal motion generation framework that decouples the conditioned generation learning process of control signals at different granularities, enabling efficient and plug-and-play multi-modal motion generation training.
  • We develop ControlMM-Bench, the first publicly available multi-modal whole-body motion generation benchmark with unified representation. Extensive experiments show that ControlMM achieves competitive performance across various standard motion generation tasks.
@@
Conditioned human motion generation models have extensive downstream applications, including text-to-motion (T2M), speech-to-gesture (S2G), and music-to-dance (M2D). In text-to-motion, models (Tevet et al. 2023; Chen et al. 2023; Zhang et al. 2023b; Liu et al. 2023; Zhang et al. 2024a, 2023c; Liang et al. 2024) achieve text-controlled motion generation with semantic consistency by applying advanced generative models and aligning motion and text feature domains. For speech-to-gesture (Yi et al. 2023; Chen et al. 2024; Liu et al. 2022b), many efforts focus on mapping speech to human gestures through rhythm alignment and character style learning. Additionally, numerous studies (Li et al. 2023; Tseng, Castellon, and Liu 2023; Siyao et al. 2022) design spatial and temporal coherence constraints to ensure that models learn the corresponding style and rhythm from the input music. Recently, increasing attention has been given to multi-modal human motion generation (Zhou, Wan, and Wang 2023; Ling et al. 2023; Zhang et al. 2024b; Luo et al. 2024). $M^{3}$-GPT (Luo et al. 2024) injects quantized compressed tokens of different modalities into the extended vocabulary of a large language model, achieving motion understanding and generation while ignoring the modeling of human topology prior. Motion-Verse (Zhang et al. 2024b) introduces dynamic attention scores to measure the relationships between different body parts, but it overlooks modeling the general basic static human topology, resulting in limited generalization and increased optimization difficulty. Additionally, it adopts the mixed training strategies for all conditions based on ImageBind (Girdhar et al. 2023), causing model confusion when learning conditions of varying granularity simultaneously and significant computation costs when retraining for new control signals. MCM (Ling et al. 2023) attempts to address the optimization confusion brought by mixed training based on utilizing the ControlNet (Zhang, Rao, and Agrawala 2023) architecture, but its channel-based attention design fails to capture either the static or the dynamic skeletal topology, resulting in poor generalization across generation scenarios. Moreover, MCM also only focuses on tensor movements, lacking the ability to generate whole-body motion. Compared with previous methods in Tab. 1, ControlMM achieve whole-body motion generation under varying control signals with plug-and-play control by adopting ControlMM-Attn to parallel capture the static human topology structure and domain-specific dynamic skeleton relationships, and conducting the coarse-to-fine conditioned generation training strategy.
@@
Various conditioned human motion generation datasets have been constructed in recent years. For text-to-motion, researchers have curated datasets encompassing action categories (Carreira et al. 2019; Chung et al. 2021; Gu et al. 2018; Liu et al. 2019; Trivedi, Thatipelli, and Sarvadevabhatla 2021), sequential action labels (Taheri et al. 2020; Zhang et al. 2022; Guo et al. 2020; Ionescu et al. 2013; Punnakkal et al. 2021), and arbitrary natural language descriptions (Lin et al. 2023; Guo et al. 2022; Tang et al. 2023) at various abstraction levels. Specifically, AMASS (Mahmood et al. 2019) consolidates 15 optical marker-based motion capture datasets into a comprehensive collection based on SMPL (Loper et al. 2015) representation. HumanML3D (Guo et al. 2022) extracted a high-quality subset within AMASS based on H3D format for torso-only generation, including three arbitrary natural language descriptions per motion clip from diverse annotators. For music-to-dance (Li et al. 2022; Zhuang et al. 2022), the most popular dataset is AIST++ (Li et al. 2021), which reconstructs 5 hours of dance based on SMPL format from videos, despite its significant reconstruction error, and lacks capture of hand movements. Finedance (Li et al. 2023) addresses these limitations by collecting 14.6 hours across 22 dance genres and supplementing the dataset with detailed hand modeling using the SMPL-H (Pavlakos et al. 2019) based representation. For speech-to-gesture, datasets (Liu et al. 2024a, 2022a; Yi et al. 2023; Ginosar et al. 2019) are gathered from pseudo-labeled (PGT) and motion-captured sources. Mocap datasets are generally preferred due to significant errors in monocular 3D pose estimation in PGT (Gärtner et al. 2022). Recently, BEATS2 (Liu et al. 2024a) and BEATS (Liu et al. 2022a) have emerged as the most popular benchmarks, celebrated for their diverse range of motion and extensive data volume. BEATS2, building upon BEATS, utilizes SMPL-X (Pavlakos et al. 2019) and FLAME (Kim, Kim, and Choi 2023) to achieve higher-quality unified mesh-level data. Despite these developments, there is no publicly available benchmark that supports unified representation for multi-modal whole-body motion generation.
@@
The key challenge in achieving whole-body human motion generation with multi-modal control is addressing motion distribution drift across different generation scenarios (Zhang et al. 2024b) and the efficient learning of control signals at varying granularities (Ling et al. 2023).
@@
Efficient learning of conditions at varying granularities. Different motion generation scenarios correspond to different control conditions at varying granularities. For instance, text guidance typically provides sequence-level coarse-grained semantic control, while first-perspective speech and third-perspective music focus more on frame-level low-level control (Liu et al. 2024a; Li et al. 2023). Mixed learning of all control signals within a single compressed latent space leads to inevitable modality alignment loss and fails to decouple the learning process for each granularity (Zhang, Rao, and Agrawala 2023; Ling et al. 2023), causing optimization confusion. Motivated by other vision generation paradigms in the image/video domain, including StableDiffusion (Rombach et al. 2022) and Sora (Liu et al. 2024b), decoupling the conditioned generation under different control signals and using T2X as a fundamental pretraining task can establish robust generative capabilities for subsequent multi-condition generation, resulting in more efficient and fine-grained multi-modal control generation capabilities.
@@
Stage  2 Multi-modal Low-level Control Adaptation. During the low-level control adaptation fine-tuning stage, we aim to model the correlation between various condition signals $\mathbf{H}_{c}\in\mathbb{R}^{B\times T_{c}\times D_{c}}$ and motion sequences $\mathbf{H}_{motion}\in\mathbb{R}^{B\times F_{m}\times D_{m}}$. All main branch parameters $f_{m}(\cdot)$ are frozen to maintain their coarse-grained motion generation and semantic text guidance following abilities. A copy of the main branch parameters $\hat{f_{m}}(\cdot)$ is then used to initialize the control branch, connecting them with a zero-initialized linear layer $\mathbf{W}_{p}\in\mathbb{R}^{D_{m}\times D_{m}}$ to prevent early training noise from causing collapse (Zhang, Rao, and Agrawala 2023). Then the condition signals (speech, music, or other low-level control signals) are fed into the control branch, where a position mask $\mathbf{M}_{c}\in\{0,1\}^{F_{m}}$ aligns the condition signals to the motion sequence length $F_{m}$, setting zeros for the $F_{m}-T_{c}$ missing frames in the original control signal sequence. Next, we employ position embeddings $\mathbf{H}_{p}\in\mathbb{R}^{F_{m}\times D_{m}}$ to explicitly align each frame of the motion sequence with the condition signals. Also, we randomly mask the condition signals with a probability of $10\%$ to achieve classifier-free guidance. The output of each control branch layer is directly added to the corresponding main branch layer input through the zero bridge linear, allowing new control signals to guide frame-level human motion generation.
@@
Various studies have shown that basic attention is sufficient for modeling temporal relationships (Nie et al. 2022). Therefore, we chose to use each body part as a unit $\mathbf{H}_{t}\in\mathbb{R}^{B\cdot N_{b}\times F_{m}\times D_{b}}$ and measure the temporal relationships between frames based on attention (Vaswani et al. 2017), which is $\hat{\mathbf{E}_{t}}=Softmax(\mathbf{Q}_{H_{t}}\mathbf{K}_{H_{t}}^{T}/\sqrt{D_{b}})\cdot\mathbf{V}_{H_{t}}$, where $\mathbf{Q}_{H_{t}}=\mathbf{W}^{Q_{H_{t}}}\mathbf{H}_{t}$, $\mathbf{K}_{H_{t}}=\mathbf{W}^{K_{H_{t}}}\mathbf{H}_{t}$ and $\mathbf{V}_{H_{t}}=\mathbf{W}^{V_{H_{t}}}\mathbf{H}_{t}$. Considering that external textual control signals are mostly sequential instructions in the temporal dimension, text information is also modeled here to produce the output $\mathbf{E}_{t}=Softmax(\mathbf{Q}_{\hat{{E}_{t}}}\mathbf{K}_{{H}_{text}}^{T}/\sqrt{D_{b}})\cdot\mathbf{V}_{{H}_{text}}$. Other sequential control modalities, such as speech and music, are modeled in the control branch. The final output $\mathbf{E}=\mathbf{E}_{s}+\mathbf{E}_{d}+\mathbf{E}_{t}$ of ControlMM-Attn combines the spatiotemporal representations of the human skeleton and the temporal structure of each body part.
@@
To avoid information loss from aligning different motion representations, we selected HumanML3D (Guo et al. 2022) in SMPL format for Text-to-Motion, FineDance (Li et al. 2023) in SMPL-H Rot-6D format for Music-to-Dance, and BEATS2 (Liu et al. 2024a) in SMPL-X format for Speech-to-Gesture from public datasets, which are the most representative unimodal datasets in their respective fields. To support whole-body multi-modal control human motion generation, we first processed all data into SMPL-X format. Key operations include filling in missing facial information in HumanML3D and FineDance with average expressions, converting the original FineDance data from SMPL-H Rot-6D format to axis-angle representation for direct alignment with SMPL-X parameters, which is more efficient than the official body-retargeting method with almost no alignment errors. We then pre-trained a motion encoder and a text encoder by aligning text and motion contrastively with a retrieval optimization goal (Lu et al. 2023) for our unified evaluation of the SMPL-X motion representation. For FineDance and BEATS2, which lack corresponding textual information, we generated pseudo-captions to serve as textual inputs, such as ”A dancer is performing a Jazz dance in the street style to the rhythm of the wildfire” and ”A person is giving a speech, and the content is …”.

### v2

# Adding Multimodal Controls to Whole-body Human Motion Generation

@@ Abstract
Whole-body multimodal motion generation, controlled by text, speech, or music, has numerous applications including video generation and character animation. However, employing a unified model to accomplish various generation tasks with different condition modalities presents two main challenges: motion distribution drifts across different generation scenarios and the complex optimization of mixed conditions with varying granularity. Furthermore, inconsistent motion formats in existing datasets further hinder effective multimodal motion generation. In this paper, we propose ControlMM, a unified framework to Control whole-body Multi-modal Motion generation in a plug-and-play manner. To effectively learn and transfer motion knowledge across different motion distributions, we propose ControlMM-Attn, for parallel modeling of static and dynamic human topology graphs. To handle conditions with varying granularity, ControlMM employs a coarse-to-fine training strategy, including stage-1 text-to-motion pre-training for semantic generation and stage-2 multimodal control adaptation for conditions of varying low-level granularity. To address existing benchmarks’ varying motion format limitations, we introduce ControlMM-Bench, the first publicly available multimodal whole-body human motion generation benchmark based on the unified whole-body SMPL-X format. Extensive experiments show that ControlMM achieves state-of-the-art performance across various standard motion generation tasks.  Corresponding authors.
Figure 1: We present ControlMM, a whole-body motion generation framework with plug-and-play multimodal control, which can generalize robust motion generation abilities across Text-to-Motion, Speech-to-Gesture, and Music-to-Dance.
@@ Introduction
Whole-body human motion generation with multimodal control  (Zhang et al. 2024b; Liu et al. 2024a; Li et al. 2023), which produces natural and logical human movements based on multimodal conditions, has many downstream applications, including human video generation (Hu 2024) and character animation (Zhang et al. 2023a).
Recent advancements in single-conditioned human motion generation have enabled realistic human motion generation under control signals from varying modalities, including text descriptions (Guo et al. 2022; Zhang et al. 2023c), music clips (Siyao et al. 2022; Li et al. 2023), and speech segments (Liu et al. 2024a; Chen et al. 2024). However, achieving whole-body motion generation with multimodal control in a unified model encounters several serious challenges: ❶ Motion distribution drift: The motion distribution in generated tasks corresponding to different control signals exhibits significant distribution drift. Under limited available motion data compared with video/text, current methods struggle to transfer robust motion topology knowledge across different multimodal motion generation scenarios (Ling et al. 2023; Zhang et al. 2024b). ❷ Struggling optimization under mixed conditions: Current multimodal motion generation methods mainly choose to compress control signals of varying granularity into a shared latent space to learn the multimodal conditioned generation abilities (Zhou, Wan, and Wang 2023; Zhang et al. 2024b), leading to alignment issues across different modalities and optimization confusion when learning conditions at different granularity simultaneously. ❸ Non-uniform motion format and evaluation: There are no multimodal whole-body human motion generation benchmarks with unified motion format and evaluation pipelines, impeding this field’s progress.
@@
  • Struggling optimization under mixed conditions: Current multimodal motion generation work often compresses different control signals (semantic text guidance, low-level speeches on first-person perspective and the low-level music on third-person perspective) into a latent space for mixed modeling, including transformer token embedding (Zhou, Wan, and Wang 2023) and ImageBind’s (Girdhar et al. 2023) hidden space (Zhang et al. 2024b). However, due to the significant information gaps between different modalities, such compression and mixed learning leads to notable optimization burden (Team et al. 2023). Therefore, we choose to introduce different control signals through separate control branches and conduct a coarse-to-fine progressive training strategy to cultivate generation abilities under conditions of varying granularity. This also provides our model with plug-and-play, user-friendly, and efficient training properties.
  • Non-uniform motion format and evaluation: Finally, there are no high-quality multimodal whole-body human motion generation benchmarks with unified motion representation and evaluation pipelines, which prompted us to build the first such benchmarks based on the SMPL-X (Pavlakos et al. 2019) whole-body format.
@@
To address existing benchmarks’ varying motion format limitations, we also built ControlMM-Bench, the first publicly available multimodal human motion generation benchmark based on the unified whole-body representation SMPL-X, including unified data construction and evaluation pipelines. Extensive experiments show that ControlMM achieves competitive performance across various standard motion generation tasks, including text-to-motion, speech-to-gesture, and music-to-dance, compared to state-of-the-art baselines. We also provide extensive ablation studies offering valuable insights into model design, training strategies, and scaling effects for future multimodal whole-body human motion generation foundation models.
@@
  • In this work, we decompose the motion distribution drift modeling in multimodal generation into the static human skeleton structure capture and dynamic topology relationship measurement. For the first time, our ControlMM-Attn achieves parallel modeling of such human-centric spatial properties to boost whole-body motion generation against distribution drifts under varying conditions.
  • We introduce ControlMM, a two-stage coarse-to-fine multimodal motion generation framework that decouples the conditioned generation learning process of control signals at different granularities, enabling efficient and plug-and-play multimodal motion generation training.
  • We develop ControlMM-Bench, the first publicly available multimodal whole-body motion generation benchmark with unified representation. Extensive experiments show that ControlMM achieves competitive performance across various standard motion generation tasks.
@@
Conditioned human motion generation models have extensive downstream applications, including text-to-motion (T2M), speech-to-gesture (S2G), and music-to-dance (M2D). In text-to-motion, models (Tevet et al. 2023; Chen et al. 2023; Zhang et al. 2023b; Liu et al. 2023; Zhang et al. 2024a, 2023c; Liang et al. 2024) achieve text-controlled motion generation with semantic consistency by applying advanced generative models and aligning motion and text feature domains. For speech-to-gesture (Yi et al. 2023; Chen et al. 2024; Liu et al. 2022b), many efforts focus on mapping speech to human gestures through rhythm alignment and character style learning. Additionally, numerous studies (Li et al. 2023; Tseng, Castellon, and Liu 2023; Siyao et al. 2022) design spatial and temporal coherence constraints to ensure that models learn the corresponding style and rhythm from the input music. Recently, increasing attention has been given to multimodal human motion generation (Zhou, Wan, and Wang 2023; Ling et al. 2023; Zhang et al. 2024b; Luo et al. 2024). $M^{3}$-GPT (Luo et al. 2024) injects quantized compressed tokens of different modalities into the extended vocabulary of a large language model, achieving motion understanding and generation while ignoring the modeling of human topology prior. Motion-Verse (Zhang et al. 2024b) introduces dynamic attention scores to measure the relationships between different body parts, but it overlooks modeling the general basic static human topology, resulting in limited generalization and increased optimization difficulty. Additionally, it adopts the mixed training strategies for all conditions based on ImageBind (Girdhar et al. 2023), causing model confusion when learning conditions of varying granularity simultaneously and significant computation costs when retraining for new control signals. MCM (Ling et al. 2023) attempts to address the optimization confusion brought by mixed training based on utilizing the ControlNet (Zhang, Rao, and Agrawala 2023) architecture, but its channel-based attention design fails to capture either the static or the dynamic skeletal topology, resulting in poor generalization across generation scenarios. Moreover, MCM also only focuses on tensor movements, lacking the ability to generate whole-body motion. Compared with previous methods in Tab. 1, ControlMM achieve whole-body motion generation under varying control signals with plug-and-play control by adopting ControlMM-Attn to parallel capture the static human topology structure and domain-specific dynamic skeleton relationships, and conducting the coarse-to-fine conditioned generation training strategy.
@@
Various conditioned human motion generation datasets have been constructed in recent years. For text-to-motion, researchers have curated datasets encompassing action categories (Carreira et al. 2019; Chung et al. 2021; Gu et al. 2018; Liu et al. 2019; Trivedi, Thatipelli, and Sarvadevabhatla 2021), sequential action labels (Taheri et al. 2020; Zhang et al. 2022; Guo et al. 2020; Ionescu et al. 2013; Punnakkal et al. 2021), and arbitrary natural language descriptions (Lin et al. 2023; Guo et al. 2022; Tang et al. 2023) at various abstraction levels. Specifically, AMASS (Mahmood et al. 2019) consolidates 15 optical marker-based motion capture datasets into a comprehensive collection based on SMPL (Loper et al. 2015) representation. HumanML3D (Guo et al. 2022) extracted a high-quality subset within AMASS based on H3D format for torso-only generation, including three arbitrary natural language descriptions per motion clip from diverse annotators. For music-to-dance (Li et al. 2022; Zhuang et al. 2022), the most popular dataset is AIST++ (Li et al. 2021), which reconstructs 5 hours of dance based on SMPL format from videos, despite its significant reconstruction error, and lacks capture of hand movements. Finedance (Li et al. 2023) addresses these limitations by collecting 14.6 hours across 22 dance genres and supplementing the dataset with detailed hand modeling using the SMPL-H (Pavlakos et al. 2019) based representation. For speech-to-gesture, datasets (Liu et al. 2024a, 2022a; Yi et al. 2023; Ginosar et al. 2019) are gathered from pseudo-labeled (PGT) and motion-captured sources. Mocap datasets are generally preferred due to significant errors in monocular 3D pose estimation in PGT (Gärtner et al. 2022). Recently, BEATS2 (Liu et al. 2024a) and BEATS (Liu et al. 2022a) have emerged as the most popular benchmarks, celebrated for their diverse range of motion and extensive data volume. BEATS2, building upon BEATS, utilizes SMPL-X (Pavlakos et al. 2019) and FLAME (Kim, Kim, and Choi 2023) to achieve higher-quality unified mesh-level data. Despite these developments, there is no publicly available benchmark that supports unified representation for multimodal whole-body motion generation.
@@
The key challenge in achieving whole-body human motion generation with multimodal control is addressing motion distribution drift across different generation scenarios (Zhang et al. 2024b) and the efficient learning of control signals at varying granularities (Ling et al. 2023).
@@
Efficient learning of conditions at varying granularities. Different motion generation scenarios correspond to different control conditions at varying granularities. For instance, text guidance typically provides sequence-level coarse-grained semantic control, while first-perspective speech and third-perspective music focus more on frame-level low-level control (Liu et al. 2024a; Li et al. 2023). Mixed learning of all control signals within a single compressed latent space leads to inevitable modality alignment loss and fails to decouple the learning process for each granularity (Zhang, Rao, and Agrawala 2023; Ling et al. 2023), causing optimization confusion. Motivated by other vision generation paradigms in the image/video domain, including StableDiffusion (Rombach et al. 2022) and Sora (Liu et al. 2024b), decoupling the conditioned generation under different control signals and using T2X as a fundamental pretraining task can establish robust generative capabilities for subsequent multi-condition generation, resulting in more efficient and fine-grained multimodal control generation capabilities.
@@
Stage  2 Multimodal Low-level Control Adaptation. During the low-level control adaptation fine-tuning stage, we aim to model the correlation between various condition signals $\mathbf{H}_{c}\in\mathbb{R}^{B\times T_{c}\times D_{c}}$ and motion sequences $\mathbf{H}_{motion}\in\mathbb{R}^{B\times F_{m}\times D_{m}}$. All main branch parameters $f_{m}(\cdot)$ are frozen to maintain their coarse-grained motion generation and semantic text guidance following abilities. A copy of the main branch parameters $\hat{f_{m}}(\cdot)$ is then used to initialize the control branch, connecting them with a zero-initialized linear layer $\mathbf{W}_{p}\in\mathbb{R}^{D_{m}\times D_{m}}$ to prevent early training noise from causing collapse (Zhang, Rao, and Agrawala 2023). Then the condition signals (speech, music, or other low-level control signals) are fed into the control branch, where a position mask $\mathbf{M}_{c}\in\{0,1\}^{F_{m}}$ aligns the condition signals to the motion sequence length $F_{m}$, setting zeros for the $F_{m}-T_{c}$ missing frames in the original control signal sequence. Next, we employ position embeddings $\mathbf{H}_{p}\in\mathbb{R}^{F_{m}\times D_{m}}$ to explicitly align each frame of the motion sequence with the condition signals. Also, we randomly mask the condition signals with a probability of $10\%$ to achieve classifier-free guidance. The output of each control branch layer is directly added to the corresponding main branch layer input through the zero bridge linear, allowing new control signals to guide frame-level human motion generation.
@@
Various studies have shown that basic attention is sufficient for modeling temporal relationships (Nie et al. 2022; Bian et al. 2024). Therefore, we chose to use each body part as a unit $\mathbf{H}_{t}\in\mathbb{R}^{B\cdot N_{b}\times F_{m}\times D_{b}}$ and measure the temporal relationships between frames based on attention (Vaswani et al. 2017). Considering that external textual control signals are mostly sequential instructions in the temporal dimension, text information is also modeled here to produce the output, $\hat{\mathbf{E}_{t}}=Softmax(\mathbf{Q}_{H_{t}}\cdot[\mathbf{K}_{H_{t}}^{T},\mathbf{K}_{H_{text}}^{T}]/\sqrt{D_{b}})\cdot[\mathbf{V}_{H_{t}}^{T},\mathbf{V}_{H_{text}}^{T}]$, where $\mathbf{Q}_{H_{t}}=\mathbf{W}^{Q_{H_{t}}}\mathbf{H}_{t}$, $\mathbf{K}_{H_{t}}=\mathbf{W}^{K_{H_{t}}}\mathbf{H}_{t}$, $\mathbf{K}_{H_{text}}=\mathbf{W}^{K_{H_{text}}}\mathbf{H}_{text}$, $\mathbf{V}_{H_{t}}=\mathbf{W}^{V_{H_{t}}}\mathbf{H}_{t}$, $\mathbf{V}_{H_{text}}=\mathbf{W}^{V_{H_{text}}}\mathbf{H}_{text}$, and [,][,][ , ] denotes concat operation. Other sequential control modalities, such as speech and music, are modeled in the control branch. The final output $\mathbf{E}=\mathbf{E}_{s}+\mathbf{E}_{d}+\mathbf{E}_{t}$ of ControlMM-Attn combines the spatiotemporal representations of the human skeleton and the temporal structure of each body part.
@@
To avoid information loss from aligning different motion representations, we selected HumanML3D (Guo et al. 2022) in SMPL format for Text-to-Motion, FineDance (Li et al. 2023) in SMPL-H Rot-6D format for Music-to-Dance, and BEATS2 (Liu et al. 2024a) in SMPL-X format for Speech-to-Gesture from public datasets, which are the most representative unimodal datasets in their respective fields. To support whole-body multimodal control human motion generation, we first processed all data into SMPL-X format. Key operations include filling in missing facial information in HumanML3D and FineDance with average expressions, converting the original FineDance data from SMPL-H Rot-6D format to axis-angle representation for direct alignment with SMPL-X parameters, which is more efficient than the official body-retargeting method with almost no alignment errors. We then pre-trained a motion encoder and a text encoder by aligning text and motion contrastively with a retrieval optimization goal (Lu et al. 2023) for our unified evaluation of the SMPL-X motion representation. For FineDance and BEATS2, which lack corresponding textual information, we generated pseudo-captions to serve as textual inputs, such as ”A dancer is performing a Jazz dance in the street style to the rhythm of the wildfire” and ”A person is giving a speech, and the content is …”.